OpenStudy (anonymous):

WILL MEDAL AND FANNN!!Which is the best estimate for the quotient of 404.31 ÷ 48.031? A. 8,000 B. 800 C. 80 D. 8

OpenStudy (kittiwitti1):

Depends on what they want you to round up to.

OpenStudy (kittiwitti1):

It looks like they want whole numbers. Here's a rule to remember for any digit rounding (digit x)\[x\lt5,\text{ round down}\rightarrow x\ge5,\text{ round up}\]

OpenStudy (kittiwitti1):

Well, specifically I think they want 10s and 100s.\[404.31\rightarrow\text{ Digit x is 4, round down to 400}\]\[48.031\rightarrow\text{ Digit x is 8, round up to 50}\]\[\text{So }\frac{404.31}{48.031}\rightarrow\frac{400}{50}=??\]

OpenStudy (anonymous):

I'm confused

OpenStudy (kittiwitti1):

What are you confused about?

OpenStudy (anonymous):

So 404.31/48.031 > 400/50=??

OpenStudy (anonymous):

that part

OpenStudy (kittiwitti1):

Yes.